Can ny1 plzz tel the integration of x4+1/x6+1..??? fast...

Grade:11

1 Answers

Aman Bansal
11 years ago

Dear Ranjita,

The integration gives: x5/5 + ln((x^7)/(7) + lx)

1/(x^6+1) = (1/3)(1/(x^2+1)) + (2/3-x^2/3)/(x^4-x^2+1).

The next step is to factor 

x^4-x^2+1 = (x^2-sqrt(3) x +1)(x^2+sqrt(3)+1) 

and do partial fractions on 

1/(x^4-x^2+1) 

Sorry about that square root of three.

PS. A clever way to see the factorization is to write 

x^4-x^2+1= (x^2+1)^2-3x^2

which is a difference of squares that is easily factored.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert